OpenStudy (anonymous):

If a point is selected at random in the interior of a circle, find the probability that the point is closer to the center than to the circle.

OpenStudy (anonymous):

To the circle's circumference?

OpenStudy (anonymous):

I don't know that is all the question says

OpenStudy (anonymous):

All right, so we have a circle with radius \(r\), and we know the point is closer to the center when it is within a radius of the center of \(\frac{r}{2}\). So, we compute the area of the circle to find that: \[ \pi \left(\frac{r}{2}\right)^2=\pi\frac{r^2}{4} \]We know the area of the outside part (or half) of the circle is: \[ \pi r^2-\pi\frac{r^2}{4}=\pi\left(r^2-\frac{r^2}{4}\right)=\pi r^2\left(1-\frac{1}{4}\right)=\pi r^2\frac{3}{4} \]So, we take the ratio between the areas, and we get a probability of \(\frac{1}{4}\)

OpenStudy (anonymous):

Sorry, of \(\frac{1}{3}\).

OpenStudy (accessdenied):

I do believe that the probability is calculated by: \[ \ \frac{\text{favorable outcomes}}{\text{total outcomes}} \] What you were calculating is the odds of choosing the specified point: \[ \ \frac{\text{favorable outcomes}}{\text{unfavorable outcomes}} \]

OpenStudy (anonymous):

Well, yeah, then in the case it would be \(\frac{1}{4}\) as my first post said.

OpenStudy (anonymous):

wait i don't get why you subtract and everything. can't u just divide your answer in the first equation by the original pi times r squared

OpenStudy (accessdenied):

Yes, that is the correct process. You need only calculate the area of the part closest to the center, and then take the ratio closest/total.

OpenStudy (anonymous):

^Eeyup, sorry, for some reason or another I thought you were looking for odds as I was typing, rather than probability...